# Color-superconducting quarkyonic matter

### v1

@@ I Introduction
The significant advances in observational astronomy over recent years have made it possible to obtain accurate measurements of various neutron star (NS) characteristics Antoniadis et al. (2013); Abbott et al. (2017, 2019, 2018); Miller et al. (2019); Riley et al. (2019); Cromartie et al. (2019); Miller et al. (2021); Riley et al. (2021); Fonseca et al. (2021). This provides a unique opportunity to probe the properties of extreme matter existing in the interiors of NSs, while neither the terrestrial experiments on collisions of relativistic heavy ions Iancu (2014); Shuryak (2017); Busza et al. (2018); Dietrich et al. (2020); Sorensen et al. (2024) nor numerical simulation in lattice Quantum Chromodynamics (QCD)  Bazavov et al. (2017, 2018, 2019); Ratti (2018); Borsanyi et al. (2018, 2020); Guenther (2021); Borsanyi et al. (2020); Borsányi et al. (2021) can access the baryon densities typical for these astrophysical objects. Thus, with the usual masses in the range $1.2$ - $2\leavevmode\nobreak\ \rm M_{\odot}$ and radii about $11$ - $13$ km, NSs can support the baryon densities several times the nuclear saturation density  Baym et al. (2018); Vidaña (2018); Capano et al. (2020); Kumar et al. (2024); Koehn et al. (2025).
@@
While deconfinement of quarks, arising from the dissociation of hadrons into their constituents, is an inevitable consequence of the asymptotic freedom of QCD at high densities Gross and Wilczek (1973); Politzer (1973); Glendenning (1997), the open question remains whether NSs support such high densities or, equivalently, whether deconfined quark matter exists in the NS interiors?
@@
Recent NICER measurements evidence that the radius of heavy PSR J0740+6620 with the mass above $2\rm M_{\odot}$ Salmi et al. (2024a); Dittmann et al. (2024) exceeds the radii of the objects PSR J0437-4715 Choudhury et al. (2024) and PSR J0030+0451  Vinciguerra et al. (2024); Miller et al. (2019) with the intermediate masses about $1.4\rm M_{\odot}$. Relatively small radius of the canonical mass NSs is also supported by the gravitational-wave signal GW170817 Abbott et al. (2017). These observational facts can be reconciled if the NS equation of state (EoS) is relatively soft at $2$ - $3$ nuclear saturation densities and significantly stiffens at higher densities due to a strong vector repulsion Dexheimer et al. (2021). This generates a characteristic back-bending in the mass-radius relation of NSs with masses $1.4$ - $2\rm M_{\odot}$. Such back-bending is typical in the presence of quark cores in NSs Alvarez-Castillo et al. (2019); Contrera et al. (2022); Ivanytskyi and Blaschke (2022a, b); Carlomagno et al. (2024); Blaschke et al. (2023a); Gärtlein et al. (2023); Li et al. (2024); Ivanytskyi (2025a); Gärtlein et al. (2025) but is challenging for most of the purely hadronic EoSs.
@@
According to the reported small radius of the ultralight HESS J1731-347 compact object Doroshenko et al. (2022), the mentioned back-bending extends even below one solar mass. While the compactness of the HESS J1731-347 is yet debated, the NICER data on PSR J1231-1411 Salmi et al. (2024b) suggest rather large radii of NSs with masses about $1\rm M_{\odot}$ and, consequently, stiff NS EoS at 2 nuclear saturation densities. Consistency of this second possibility with the mentioned above soft NS EoS at $1$ - $1.4\rm M_{\odot}$ can be provided by a strong first-order phase transition occurring at those NS masses. Each of the described scenarios for light mass NSs is consistent with the existence of quark matter in their interiors. The nature of the Black Widow Pulsar J0952-0607 with the mass $2.35\pm 0.17\leavevmode\nobreak\ {\rm M}_{\odot}$ Romani et al. (2022) also can be explained as an NS with color-superconducting (CS) quark core. Very recent results of physics-informed Bayesian analysis of observational data also provide strong statistical evidence in favor of the existence of quark cores in NSs compared to purely hadronic NS scenarios Ayriyan et al. (2025). In addition, as recently demonstrated Sagun et al. (2023), hybrid stars with paired quark matter in their cores show good agreement with observational data on the thermal evolution of NSs, including HESS J1731-347 Doroshenko et al. (2022).
@@
If NSs indeed contain quark matter, they provide a unique opportunity to probe the properties of the transition between quark and hadronic matter, which is of fundamental importance for understanding the QCD phase diagram Wilczek (1999); Rischke (2004); Stephanov (2006); Fukushima and Hatsuda (2011); Sorensen et al. (2024); Murgana and Ruggieri (2025); Abuali et al. (2025); Li et al. (2025); Schmidt (2025); Bonnet et al. (2025). In the case of a sharp transition between hadronic and quark degrees of freedom, the phase transition is of the first order, characterized by a discontinuous jump in density. Usually, such a phase transition is modeled by the Maxwell construction Baym et al. (2018). This is equivalent to the Gibbs criterion of phase equilibrium under the condition of baryon charge conservation, while the conservation of electric charge is loosened, which leads to a discontinuity of the corresponding chemical potential. In the case of two or more conserved charges, as when the conservation of electric charge is accounted for along with the conservation of baryon charge, the Gibbs criterion leads to a continuous behavior of the density across the quark-hadron phase transition. This corresponds to the Glendenning construction of the phase transition Glendenning (1992), which is the limiting case of the phase coexistence with the zero surface tension between the phases. A more realistic description requires accounting for spatial inhomogeneities, the so-called pasta phases (see, e.g., Ref. Maslov et al. (2019)), which are mimicked by various interpolation schemes Ayriyan et al. (2018, 2021); Ivanytskyi and Blaschke (2022c).
@@
Answering the question about the details of the quark-hadron transition and its relevance for NSs requires a unified quark-hadron approach. The picture of quarkyonic (QY) matter provides a physically transparent framework for this approach McLerran and Pisarski (2007); Hidaka et al. (2008). Its main phenomenological consequence is a smooth character of the deconfinement phase transition that leads to a pronounced peak of the speed of sound McLerran and Reddy (2019). This feature has been utilized in modeling the NS interiors McLerran and Reddy (2019); Zhao and Lattimer (2020); Zhang and Mann (2021); Park et al. (2021); Koch and Vovchenko (2023); Cao (2022); Pang et al. (2024); Poberezhnyuk et al. (2023); Duarte et al. (2023); Gao and Harada (2025); Folias and Moustakidis (2025); Dey et al. (2025a); Gao and Harada (2025); Fujimoto et al. (2024a, 2025); Dey et al. (2025b) reaching the high masses consistent with observational constraints Antoniadis et al. (2013); Fonseca et al. (2021). Recently, the QY picture has also been applied for modeling cold QCD matter at vanishing baryon densities and large isospin asymmetries Ivanytskyi (2025b).
QY matter might give rise to distinct dynamical signatures, for instance, in binary NS mergers. Thus, in contrast to a sharp deconfinement phase transition that generates resonant tidal excitations of interfacial i-modes during the binary NS coalescence Counsell et al. (2025), QY matter is not expected to produce this effect. The smooth onset of quarks does not create a distinct interface, thereby making its detection considerably more challenging.
@@
While resembling the scenario of quark-hadron continuity Schäfer and Wilczek (1999); Baym et al. (2018); Kojo (2021); Fujimoto et al. (2024b) and being supported by holographic models Kovensky and Schmitt (2020); Yang and Yuan (2022); Chen et al. (2020), the QY picture is based on the assumption that quarks remain bound to hadrons up to the quark chemical potentials $\mu_{q}$ being parametrically large compared to the QCD energy scale $\Lambda$. The ’t Hooft limit of a large number of colors $N_{c}$ and the vanishing QCD coupling as $g_{\rm QCD}\propto 1/N_{c}$ allow for a simple illustration of this picture. The confining interactions are suppressed when the Debye mass exceeds the QCD energy scale. At zero temperature typical for the NS interiors Fukushima and Hatsuda (2011); Kurkela et al. (2010); Baym et al. (2018), the Debye mass scales as $g_{\rm QCD}^{2}N_{c}\mu_{q}^{2}$. With this, we conclude that the confining interactions remain important up to $\mu_{q}\simeq\sqrt{N_{c}}\Lambda$ [^1] In other words, the deconfinement chemical potential being parametrically large compared to the QCD scale assumes the scaling $\mu_{B}/\Lambda\simeq\sqrt{N_{c}}$, which is large in the ’t Hooft limit.. This binds quarks from the vicinity of their Fermi sphere of size $k_{h}$ to hadrons. The momentum states resided deeply inside of the Fermi sea and having momenta smaller than $k_{q}$ do not experience the confinement since the Pauli blocking among them disables the confining scatterings McLerran and Pisarski (2007); Hidaka et al. (2008). Thus, the QY picture suggests that quarks with $k_{q}\leq|{\bf k}|\leq k_{h}$ are confined to hadrons, while the states with $|{\bf k}|<k_{q}$ remain effectively unbound. At small densities, $k_{q}$ vanishes so there are no unbound quarks, while at high densities $k_{q}$ approaches $k_{h}$ and hadrons disappear due to their dissociation to the constituent quarks.
@@
The described QY picture is similar to the phenomenon of Cooper pairing, which also mostly affects the high-momentum states in the vicinity of the Fermi sphere Cooper (1956); Bardeen et al. (1957). Accounting for this and for the possible interpretation of baryons as color singlet bound states of quarks and diquarks  Praschifka et al. (1987); Cahill (1989); Cahill et al. (1989); Burden et al. (1989); Reinhardt (1990); Zuckert et al. (1997); Wang et al. (2011); Blanquier (2011), we conclude that diquarks should also play a role in QY matter where baryons are constituted by the quark states from the vicinity of the Fermi sphere. Bose-Einstein condensation of diquarks corresponds to the formation of CS Son (1999); Alford et al. (1999); Abuki (2003); Yuan et al. (2024), which plays an important role in the phenomenology of NSs Buballa (2005); Baym et al. (2018). It is worth mentioning that since superconductivity is inevitable for any sufficiently dense fermionic system with an arbitrarily weak attractive interaction Cooper (1956), the CS should necessarily manifest itself in QY matter. However, to the best of our knowledge, CS has never been accounted for within the framework of QY matter. The present paper fills this gap, introducing a model of CS quarkyonic (CSQY) matter in the case of two-flavor CS (2CS) Buballa (2005); Baym et al. (2018). We also model NSs with CSQY cores and show that this scenario provides a good agreement with the observational data.
@@ II Quarks in Color-Superconducting Quarkyonic matter
In the presence of the 2SC phase, only two color states of quarks are paired, while the third one is not Buballa (2005). Conventionally, we define the red and green color states to be paired, while the blue color state is unpaired. The phenomenological picture of CSQY matter consistent with this convention and with the notations introduced in Sec. I is shown by Fig. 1. All the color states of quarks are labeled by the subscript index $c=(r,g,b)$. Thus, below we quantify the effects of the 2SC phase by the color-dependent pairing gap $\Delta_{c{\bf k}}=(\Delta_{\bf k},\Delta_{\bf k},0)$. The momentum dependence of the gap $\Delta_{\bf k}$ is motivated by the running of the QCD coupling Deur et al. (2016) and is introduced similarly to the nonlocal Nambu–Jona–Lasinio (NJL) model Blaschke et al. (2023b); Contrera et al. (2022); Carlomagno et al. (2024); Ivanytskyi (2025a).
The color dependence of the 2SC matter pairing gap leads to an inequivalence among quarks of different colors, resulting in distinct Fermi momenta $k_{fc}$ that generally do not coincide. Hereafter, the index $f=(u,d)$ labels quark flavor states with masses $m_{f}$. In comparison to the chiral quark models, in which the medium-dependent quark masses are defined within a self-consistent mean-field approximation (see e.g. Refs. Buballa (2005); Ivanytskyi and Blaschke (2022a); Ivanytskyi (2025a)), for the sake of simplicity, we treat the quark masses as constant parameters of the model. We also allow them to deviate from the current quark masses from the Review of Particle Physics Workman et al. (2022) and assume that they partially account for the complexity of quark in-medium interactions. As is shown below, the Fermi energy of a given flavor-color state of quarks is $\epsilon_{fc}-\Delta_{fc}$, where $\Delta_{fc}$ denotes the pairing gap defined at the quark Fermi momentum and $\epsilon_{fc}=\sqrt{k_{fc}^{2}+m_{f}^{2}}$.
The single-particle distribution function of unpaired blue quarks is given by the Fermi-Dirac distribution controlled by the corresponding Fermi momentum. The single-particle distribution of paired quarks can be derived within the Nambu-Gorkov formalism Nambu (1960); Gor’kov (1958). However, this formalism treats quarks of all momenta as unbound quasiparticles, while the striking element of the QY picture is that the momentum states above the Fermi sphere are bound to baryons McLerran and Reddy (2019). This circumstance can be accounted for by cutting off from the Nambu-Gorkov single-particle distribution the momentum states above $k_{fc}$.
To define the single-particle distribution function of quarks at a given color-flavor state in CSQY matter, we start with the corresponding Nambu-Gorkov number density Buballa (2005); Baym et al. (2018); Ivanytskyi and Blaschke (2022a); Ivanytskyi (2025a)
@@
At a vanishing pairing gap, the Nambu-Gorkov distribution function (4) reduces to the zero-temperature Fermi-Dirac distribution $\theta(\epsilon_{fc}-\epsilon_{f{\bf k}})$. At a finite pairing gap below the Fermi surface $\tilde{f}_{fc{\bf k}}<\theta(\epsilon_{fc}-\epsilon_{f{\bf k}})=1$ and $\tilde{f}_{fc{\bf k}}>\theta(\epsilon_{fc}-\epsilon_{f{\bf k}})=0$ above it. This manifests in the redistribution of the quark momentum states mentioned above (see Fig. 2). Cutting off the states above the Fermi sphere produces the single-particle distribution of quarks in CSQY matter. Thus, we define
@@
gets reduced compared to the corresponding density of normal QY matter. This effect is reflected in the CS induced stiffening of the QY matter [^2] The review Baym et al. (2018) provides an excellent discussion of stiffness of quark-hadron EoSs., which is discussed in Sec. IV.

### v2

@@ I Introduction
The significant advances in observational astronomy over recent years have made it possible to obtain accurate measurements of various neutron star (NS) characteristics Antoniadis:2013pzd; LIGOScientific:2017vwq; LIGOScientific:2018hze; LIGOScientific:2018cki; Miller:2019cac; Riley:2019yda; NANOGrav:2019jur; Miller:2021qha; Riley:2021pdl; Fonseca:2021wxt. This provides a unique opportunity to probe the properties of extreme matter existing in the interiors of NSs, while neither the terrestrial experiments on collisions of relativistic heavy ions Iancu:2012xa; Shuryak:2014zxa; Busza:2018rrf; Dietrich:2020efo; Sorensen:2023zkk nor numerical simulation in lattice Quantum Chromodynamics (QCD) Bazavov:2017dus; Bazavov:2017dsy; HotQCD:2018pds; Ratti:2018ksb; Borsanyi:2018grb; Borsanyi:2020fev; Guenther:2020jwe; Borsanyi:2020fev; Borsanyi:2021sxv can access the baryon densities typical for these astrophysical objects. Thus, with the usual masses in the range $1.2$ - $2\penalty 10000\ \rm M_{\odot}$ and radii about $11$ - $13$ km, NSs can support the baryon densities several times the nuclear saturation density Baym:2017whm; Vidana:2018lqp; Capano:2019eae; MUSES:2023hyz; Koehn:2024set.
@@
While deconfinement of quarks, arising from the dissociation of hadrons into their constituents, is an inevitable consequence of the asymptotic freedom of QCD at high densities Gross:1973id; Politzer:1973fx; Glendenning:1997wn, the open question remains whether NSs support such high densities or, equivalently, whether deconfined quark matter exists in the NS interiors?
@@
Recent NICER measurements evidence that the radius of heavy PSR J0740+6620 with the mass above $2\rm M_{\odot}$ Salmi:2024aum; Dittmann:2024mbo exceeds the radii of the objects PSR J0437-4715 Choudhury:2024xbk, PSR J0030+0451 Vinciguerra:2023qxq; Miller:2019cac and PSR J0614-3329 Mauviard:2025dmd with the intermediate masses about $1.4\rm M_{\odot}$. Relatively small radius of the canonical mass NSs is also supported by the gravitational-wave signal GW170817 LIGOScientific:2017vwq. These observational facts can be reconciled if the NS equation of state (EoS) is relatively soft at $2$ - $3$ nuclear saturation densities and significantly stiffens at higher densities due to a strong vector repulsion Dexheimer:2020rlp. This generates a characteristic back-bending in the mass-radius relation of NSs with masses $1.4$ - $2\rm M_{\odot}$. Such back-bending is typical in the presence of quark cores in NSs Alvarez-Castillo:2018pve; Contrera:2022tqh; Ivanytskyi:2022oxv; Ivanytskyi:2022bjc; Carlomagno:2023nrc; Blaschke:2023; Gartlein:2023vif; Li:2024lmd; Ivanytskyi:2024zip; Gartlein:2024cbj but is challenging for most of the purely hadronic EoSs.
@@
According to the reported small radius of the ultralight HESS J1731-347 compact object Doroshenko2022, the mentioned back-bending extends even below one solar mass. While the compactness of the HESS J1731-347 is yet debated, the NICER data on PSR J1231-1411 Salmi:2024bss suggest rather large radii of NSs with masses about $1\rm M_{\odot}$ and, consequently, stiff NS EoS at 2 nuclear saturation densities. Consistency of this second possibility with the mentioned above soft NS EoS at $1$ - $1.4\rm M_{\odot}$ can be provided by a strong first-order phase transition occurring at those NS masses or even below them. A strong first-order phase transition at small NS mass can also explain the reported small radius of the HESS J1731-347 object making it consistent with the existence of quark matter in the interiors of all NSs Sagun:2023rzp. The nature of the Black Widow Pulsar J0952-0607 with the mass $2.35\pm 0.17\penalty 10000\ {\rm M}4_{\odot}$ Romani:2022jhd also can be explained as an NS with color-superconducting (CS) quark core. Very recent results of physics-informed Bayesian analysis of observational data also provide strong statistical evidence in favor of the existence of quark cores in NSs compared to purely hadronic NS scenarios Ayriyan:2025rub. In addition, as recently demonstrated Sagun:2023rzp, hybrid stars with paired quark matter in their cores show good agreement with observational data on the thermal evolution of NSs, including HESS J1731-347 Doroshenko2022.
@@
If NSs indeed contain quark matter, they provide a unique opportunity to probe the properties of the transition between quark and hadronic matter, which is of fundamental importance for understanding the QCD phase diagram Wilczek:1999ym; Rischke:2003mt; Stephanov:2006zvm; Fukushima:2010bq; Sorensen:2023zkk; Murgana:2025wsh; Abuali:2025tbd; Li:2025ugv; Schmidt:2025ppy; INDRA:2025htq. In the case of a sharp transition between hadronic and quark degrees of freedom, the phase transition is of the first order, characterized by a discontinuous jump in density. Usually, such a phase transition is modeled by the Maxwell construction Baym:2017whm. This is equivalent to the Gibbs criterion of phase equilibrium under the condition of baryon charge conservation, while the conservation of electric charge is loosened, which leads to a discontinuity of the corresponding chemical potential. In the case of two or more conserved charges, as when the conservation of electric charge is accounted for along with the conservation of baryon charge, the Gibbs criterion leads to a continuous behavior of the density across the quark-hadron phase transition. This corresponds to the Glendenning construction of the phase transition Glendenning:1992vb, which is the limiting case of the phase coexistence with the zero surface tension between the phases. A more realistic description requires accounting for spatial inhomogeneities, the so-called pasta phases (see, e.g., Ref. Maslov:2018ghi), which are mimicked by various interpolation schemes Ayriyan:2017nby; Ayriyan:2021prr; Ivanytskyi:2022wln; Gao:2025vdc.
@@
Answering the question about the details of the quark-hadron transition and its relevance for NSs requires a unified quark-hadron approach. The picture of quarkyonic (QY) matter provides a physically transparent framework for this approach McLerran:2007qj; Hidaka:2008yy. Its main phenomenological consequence is a smooth character of the deconfinement phase transition that leads to a pronounced peak of the speed of sound McLerran:2018hbz. This feature has been utilized in modeling the NS interiors McLerran:2018hbz; Zhao:2020dvu; Zhang:2020jmb; Park:2021hqb; Koch:2022act; Cao:2022inx; Pang:2023dqj; Poberezhnyuk:2023rct; Duarte:2023cki; Gao:2024jlp; Folias:2024upz; Dey:2024lco; Gao:2024jlp; Fujimoto:2024doc; Fujimoto:2025sxx; Dey:2025jbm; Bluhm:2024uhj reaching the high masses consistent with observational constraints Antoniadis:2013pzd; Fonseca:2021wxt. Recently, the QY picture has also been applied for modeling cold QCD matter at vanishing baryon densities and large isospin asymmetries Ivanytskyi:2025cnn.
QY matter might give rise to distinct dynamical signatures, for instance, in binary NS mergers. While a sharp deconfinement phase transition generates resonant tidal excitations of interfacial i-modes during the binary NS coalescence Counsell:2025hcv, QY matter might produce a different effect. The smooth onset of quarks does not create a distinct interface, thereby making its gravitational wave imprint very different from the one of a first-order quark-hadron transition with a systematic decrease of the dominant postmerger frequency Bauswein:2018bma. This effect is caused by vanishing incompressibility of the mixed phase at a first-order phase transition and is not observed in the case of a smooth mixed phase modeled within the Glendenning construction Prakash:2021wpz or with an interpolation scheme Kedia:2022nns. Relative stiffening of a smooth mixed phase region compared to the sharp one leads to prolonging the postmerger stage but does not significantly modify the dominant postmerger frequency compared to the purely hadronic scenario. Based on this we can expect that significant stiffness of the QY matter right after the onset of quark matter (see e.g. McLerran:2018hbz) can prolong the postmerger phase even more and increase its dominant frequency.
@@
While resembling the scenario of quark-hadron continuity Schafer:1998ef; Baym:2017whm; Kojo:2021ugu; Fujimoto:2023mzy and being supported by holographic models Kovensky:2020xif; Yang:2020hun; Chen:2019rez, the QY picture is based on the assumption that quarks remain bound to hadrons up to the quark chemical potentials $\mu_{q}$ being parametrically large compared to the QCD energy scale $\Lambda$. The ’t Hooft limit of a large number of colors $N_{c}$ and the vanishing QCD coupling as $g_{\rm QCD}\propto 1/N_{c}$ allow for a simple illustration of this picture. The confining interactions are suppressed when the Debye mass exceeds the QCD energy scale. At zero temperature typical for the NS interiors Fukushima:2010bq; Kurkela:2009gj; Baym:2017whm, the Debye mass scales as $g_{\rm QCD}^{2}N_{c}\mu_{q}^{2}$. With this, we conclude that the confining interactions remain important up to $\mu_{q}\simeq\sqrt{N_{c}}\Lambda$ [^1] In other words, the deconfinement chemical potential being parametrically large compared to the QCD scale assumes the scaling $\mu_{B}/\Lambda\simeq\sqrt{N_{c}}$, which is large in the ’t Hooft limit.. This binds quarks from the vicinity of their Fermi sphere of size $k_{h}$ to hadrons. The momentum states resided deeply inside of the Fermi sea and having momenta smaller than $k_{q}$ do not experience the confinement since the Pauli blocking among them disables the confining scatterings McLerran:2007qj; Hidaka:2008yy. Thus, the QY picture suggests that quarks with $k_{q}\leq|{\bf k}|\leq k_{h}$ are confined to hadrons, while the states with $|{\bf k}|<k_{q}$ remain effectively unbound. At small densities, $k_{q}$ vanishes so there are no unbound quarks, while at high densities $k_{q}$ approaches $k_{h}$ and hadrons disappear due to their dissociation to the constituent quarks.
@@
The described QY picture is similar to the phenomenon of Cooper pairing, which also mostly affects the high-momentum states in the vicinity of the Fermi sphere Cooper:1956zz; Bardeen:1957mv. Accounting for this and for the possible interpretation of baryons as color singlet bound states of quarks and diquarks Praschifka:1986nf; Cahill:1988zi; Cahill:1988bh; Burden:1988dt; Reinhardt:1989rw; Zuckert:1996nu; Wang:2010iu; Blanquier:2011zz, we conclude that diquarks should also play a role in QY matter where baryons are constituted by the quark states from the vicinity of the Fermi sphere. Bose-Einstein condensation of diquarks corresponds to the formation of CS Son:1998uk; Alford:1999pa; Abuki:2003ut; Yuan:2024ajk, which plays an important role in the phenomenology of NSs Buballa:2003qv; Baym:2017whm. It is worth mentioning that since superconductivity is inevitable for any sufficiently dense fermionic system with an arbitrarily weak attractive interaction Cooper:1956zz, the CS should necessarily manifest itself in QY matter. However, to the best of our knowledge, CS has never been accounted for within the framework of QY matter. The present paper fills this gap, introducing a model of CS quarkyonic (CSQY) matter in the case of two-flavor CS (2CS) Buballa:2003qv; Baym:2017whm. We also model NSs with CSQY cores and show that this scenario provides a good agreement with the observational data.
@@ II Quarks in Color-Superconducting Quarkyonic matter
In the presence of the 2SC phase, only two color states of quarks are paired, while the third one is not Buballa:2003qv. Conventionally, we define the red and green color states to be paired, while the blue color state is unpaired. The phenomenological picture of CSQY matter consistent with this convention and with the notations introduced in Sec. I is shown in Fig. 1. All the color states of quarks are labeled by the subscript index $c=(r,g,b)$. Thus, below we quantify the effects of the 2SC phase by the color-dependent pairing gap $\Delta_{c{\bf k}}=(\Delta_{\bf k},\Delta_{\bf k},0)$. The momentum dependence of the gap $\Delta_{\bf k}$ is motivated by the running of the QCD coupling Deur:2016tte and is introduced similarly to the nonlocal Nambu–Jona–Lasinio (NJL) model Blaschke:2022egm; Contrera:2022tqh; Carlomagno:2023nrc; Ivanytskyi:2024zip.
The color dependence of the 2SC matter pairing gap leads to an inequivalence among quarks of different colors, resulting in distinct Fermi momenta $k_{fc}$ that generally do not coincide [^2] The mismatch of quark momenta in the presence of CS is discussed in the review Schmitt:2025cqi.. Hereafter, the index $f=(u,d)$ labels quark flavor states with masses $m_{f}$. In comparison to the chiral quark models, in which the medium-dependent quark masses are defined within a self-consistent mean-field approximation (see e.g. Refs. Buballa:2003qv; Ivanytskyi:2022oxv; Ivanytskyi:2024zip), for the sake of simplicity, we treat the quark masses as constant parameters of the model. We also allow them to deviate from the current quark masses from the Review of Particle Physics ParticleDataGroup:2022pth and assume that they partially account for the complexity of quark in-medium interactions. As is shown below, the Fermi energy of a given flavor-color state of quarks is $\epsilon_{fc}-\Delta_{fc}$, where $\Delta_{fc}$ denotes the pairing gap defined at the quark Fermi momentum and $\epsilon_{fc}=\sqrt{k_{fc}^{2}+m_{f}^{2}}$.
The single-particle distribution function of unpaired blue quarks is given by the Fermi-Dirac distribution controlled by the corresponding Fermi momentum. The single-particle distribution of paired quarks can be derived within the Nambu-Gorkov formalism Nambu:1960tm; Gorkov:1958. However, this formalism treats quarks of all momenta as unbound quasiparticles, while the striking element of the QY picture is that the momentum states above the Fermi sphere are bound to baryons McLerran:2018hbz. This circumstance can be accounted for by cutting off from the Nambu-Gorkov single-particle distribution the momentum states above $k_{fc}$.
To define the single-particle distribution function of quarks at a given color-flavor state in CSQY matter, we start with the corresponding Nambu-Gorkov number density Buballa:2003qv; Baym:2017whm; Ivanytskyi:2022oxv; Ivanytskyi:2024zip
@@
At a vanishing pairing gap, the Nambu-Gorkov distribution function (4) reduces to the zero-temperature Fermi-Dirac distribution $\theta(\epsilon_{fc}-\epsilon_{f{\bf k}})$. At a finite pairing gap below the Fermi surface $\tilde{f}_{fc{\bf k}}<\theta(\epsilon_{fc}-\epsilon_{f{\bf k}})=1$ and $\tilde{f}_{fc{\bf k}}>\theta(\epsilon_{fc}-\epsilon_{f{\bf k}})=0$ above it. This manifests in the redistribution of the quark momentum states mentioned above (see Fig. 2). According to the QY picture, the momentum states above the Fermi sphere (orange area in Fig. 2) are no longer part of the single particle spectrum of free quarks but redistribute to the spectrum of baryons. Therefore, these states should be cut off from $\tilde{f}_{fc{\bf k}}$ and consistently accounted for through the distribution function of baryons. This produces the single-particle distribution of quarks in CSQY matter. Thus, we define
@@
gets reduced compared to the corresponding density of normal QY matter. This effect is reflected in the CS induced stiffening of the QY matter [^3] The review Baym:2017whm provides an excellent discussion of stiffness of quark-hadron EoSs., which is discussed in Sec. IV.
@@
The described scheme of accounting for CS in QY matter by modifying the Nambu-Gorkov distribution of paired quarks suggests a reasonable compromise between the physical transparency of QY matter and the technical complexity of its complete microscopic theory in the presence of CS. Such a theory can be constructed similarly to the QY picture of isospin QCD Ivanytskyi:2025cnn. This requires a self-consistent determination of the momentum shell of bound quarks and a dynamical evaluation of the spectral properties of nucleons, which are considered as bound states of quarks in the mentioned momentum shell.